OpenStudy (anonymous):

write square root of y^2+4 in terms of theta (X) when y=2tanX

OpenStudy (anonymous):

\[\sqrt{(2\tan(x))^2+4}\] \[\sqrt{4\tan^2(x)+4}\] \[2\sqrt{\tan^2(x)+1}\] \[2\sqrt{\sec^2(x)}\] \[2 \sec(x)\] or i guess \[2|\sec(x)|\] but you usually assume x is such that sec(x) > 0

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

yw